OpenStudy (anonymous):

Rewrite the logarithmic expression as a single logarithm with a coefficient of 1. lnx + ln(x+1)

OpenStudy (anonymous):

ln is spelled log below: \[\log (x)+\log (x+1)=\log (x (x+1)) \]